The pressure was exerted with a piston on molted naphthalene (C10H18) at 95 oC. The vapour pressure of naphthalene under 1.0 bar is 2.0 kPa and its density is 0.962 g cm-3, calculate the normal boiling point.


1
Expert's answer
2021-01-20T03:15:43-0500

Standard vaporization enthalpy of naphthalene is 53.4 kJ/mol (table data). Let's assume enthalpy does not change with temperature.


Normal boiling point means vapour pressure is equal to the standard atmospheric pressure (1 atm or 1.01325 bar or 101325 Pa)


ln(p1/p2) = "\Delta"vapH/R*(1/T2-1/T1)


p2 is 2kPa and T2 is 95⁰C or 368.15K


ln(101325/2000) = 53400/8.314*(1/368.15-1/Tn.b.p.)


Tn.b.p. = 475.02K or 201.87⁰C